# LLMs cannot find reasoning errors, but can correct them!

### v2

@@ 4 Backtracking
2.
We then determine the mistake location in this trace using a reward model.
3.
If there are no mistakes, we move onto the next trace. If there is a mistake (e.g. at Thought 4 in the example trace in Table 1),we prompt the model again for the same step but at temperature , generating 8 outputs. We use same prompt and the partial trace containing all steps up to but not including the mistake step (e.g. up to Thought 3, prompting for Thought 4).
4.
From the 8 outputs, we filter out any options that match what was previously identified as a mistake. From the remaining outputs, we select one with the highest log-probability.
5.
@@
Our backtracking method provides several benefits over existing self-correction methods:
Unlike Shinn et al. (2023), Kim et al. (2023), etc., our approach does not depend on oracle knowledge of the answer. Instead, it relies on information (for example from trained a reward model) about logical errors, which can be determined on a step-by-step basis using a reward model. Logical errors can occur in correct traces, or not occur in incorrect traces888Having no logical errors in incorrect traces is much rarer but does exist, for example when the answer is correct but is not captured by exact match, or if the original question is faulty and has multiple possible answers..
Compared to approaches that require re-generating the entire trace, backtracking reduces computational cost by reusing previous steps that are known to be logically sound.
Backtracking improves on the quality of the intermediate steps directly, which can be useful in scenarios that require correct steps (e.g. generating solutions to math questions), and also generally improves interpretability.
Backtracking with mistake location information from a reward model can be construed as a lightweight RL method. However, unlike conventional deep reinforcement learning:
Backtracking with a reward model does not does not require any training of the original generator model. Once the reward model is trained, it can be used for backtracking with any LLM as the generator, and can also be updated independently of the generator LM. This can be especially helpful when LLMs are frequently updated to new checkpoints.
Backtracking only requires training of a small reward model. Compared to methods that require training of the generator model, backtracking is far more efficient in terms of computing resources and available data.
The process of backtracking is more interpretable than updating the weights of the generator model directly, as is required for many deep RL methods. It clearly pinpoints the location at which an error occurs, which can help the debugging process and allow faster development and iterations of models.
4.1 Backtracking with gold mistake location
As an initial experiment, we use labels from BIG-Bench Mistake to test if an LLM is able to correct logical errors using backtracking, independent of its inherent ability to identify these errors or any other reward model.
For example, if the mistake location is in step 4, we use backtracking to regenerate that step and continue the rest of the chain. If the mistake location is that there are no logical mistakes, we do not backtrack and use the original result.
| With mistake location | With random location |
Avg. num. of steps |
@@ 4.1 Backtracking with gold mistake location
| Word sorting | -11.11 | +23.53 | -15.56 | +11.76 |
11.7 |
| Tracking shuffled objects | -6.67 | +43.92 | -6.67 | +20.39 |
5.4 |
| Logical deduction | -11.43 | +36.86 | -13.33 | +21.57 |
8.3 |
| Multistep arithmetic | -0.00 | +18.04 | -8.89 | +10.59 |
5.0 |
| Dyck languages | -6.82 | +18.06 | -15.91 | +5.16 |
24.5 |
4.1.1 Results
The results are shown in Table 6. To show that performance increases are not due to randomly resampling outputs, we compare our results to a random baseline, where a mistake location999As described above, the mistake location can be either the number representing the step, or that there are no mistakes. If there are no mistakes, we do not use backtracking and simply use the original trace. is randomly selected for each trace and we perform backtracking based on the random location.
Note that Table 6 separates results into numbers for the correct set and the incorrect set, referring to whether the original trace was correct or not. This gives a clearer picture than the overall accuracy, which would be skewed by the proportion of traces that were originally correct (15%) and incorrect (85%).
Scores represent the absolute differences in accuracy. We perform backtracking on both correct and incorrect traces, as long as there is a mistake in one of the steps.
@@ 4.1.1 Results
refers to differences in accuracy on the set of traces whose original answer was correct. Note that we take losses here because, despite the correct answer, there is a logical mistake in one of the steps. Therefore, the answer may change to an incorrect one when we backtrack.
@@
is the same but for incorrect traces, so the answers may have been corrected, hence increasing accuracy.
@@
For example, for the word sorting task, 11.11% of traces that were originally correct became incorrect, while 23.53% of traces that were originally incorrect became correct.
4.1.2 Discussion
@@
The scores show that the gains from correcting incorrect traces are larger than losses from changing originally correct answers. Additionally, while the random baseline also obtained improvements, they are considerably smaller than if the true mistake location was used. Note that tasks involving fewer steps are more likely to improve performance in the random baseline, as the true mistake location is more likely to be identified.
@@
While our numbers do show that our gains are higher than our losses, it should be noted that changes in the overall accuracy depends on the original accuracy achieved on the task. For example, if the original accuracy on the tracking shuffled objects task was 50%, the new accuracy would be 68.6%. On the other hand, if the accuracy was 99%, the new accuracy would drop to 92.8%. As our dataset is highly skewed and only contains 45 correct traces per task, we leave to future work to assess the effectiveness of backtracking in a more comprehensive way.
4.2 Backtracking with a simulated reward model
We show in subsection 4.1 that backtracking can be used to correct CoT traces using gold mistake location labels. To explore what level of accuracy reward model is needed when gold labels are not available, we use backtracking with simulated reward models, designed to produce labels at different levels of accuracy. We use accuracy to refer to the accuracy of the simulated reward model at identifying mistake locations.
@@ 4.2 Backtracking with a simulated reward model
For a given reward model at % accuracy, we use the mistake location from BIG-Bench Mistake % of the time. For the remaining %, we sample a mistake location randomly. To mimic the behaviour of a typical classifier, mistake locations are sampled to match the distribution found in the dataset. We also ensure that the sampled location does not match the correct location.
4.2.1 Results
@@
Results are shown in Figure 2. We can see that the losses in accuracy begins to plateau at 65%. In fact, for most tasks, accuracy is already larger than accuracy at around 60-70% accuracy. This demonstrates that while higher accuracies produce better results, backtracking is still effective even without gold standard mistake location labels.
4.3 Reward Modeling
@@
We perform a preliminary investigation of if mistake-finding can benefit from a dedicated reward model and if learning to find mistakes in a set of tasks can transfer to finding mistakes in out-of-distribution tasks. We fine-tuned a PaLM 2-XS-Otter model based on our available data for 20k steps and choose the checkpoint with the best validation results. We hold out one task for evaluation while training the reward model on the other 4 tasks.
| Held-out task | Difference |
| Word sorting | +11.66 |
| Tracking shuffled objects | +19.67 |
| Logical deduction | -0.67 |
| Multi-step arithmetic | +4.00 |
| Dyck languages | +22.59 |
@@
Note the reward model we train is significantly smaller than our inference model. We show the relative improvements and losses in Table 7 vs. a zero-shot baseline on PaLM 2-L-Unicorn. We see gains for 4 out of 5 of the tasks. This provides initial indication that it maybe possible to train separate reward model classifiers to assist in backtracking and that these reward models do not have to be large. Further, a reward model can work on mistakes that are out-of-distribution. However, we believe more data may be necessary to improve results across the board on all tasks. We leave the collection of this larger dataset and a more rigorous investigation of the trade-offs of model size vs. performance of the reward model to future work.
@@
We also leave for future work the effect of backtracking iteratively with a reward model: for example, the generator model may make another mistake after backtracking for the first time, which can then be identified and corrected again.
@@ 5 Related work
Datasets
To our knowledge, the only publicly available dataset containing mistake annotations in LLM outputs is PRM800K (Lightman et al., 2023), which is a dataset of solutions to Olympiad-level math questions. Our dataset BIG-Bench Mistake covers a wider range of tasks to explore the reasoning capabilities of LLMs more thoroughly. Additionally, the generator LLM used in PRM800K has been fine-tuned on 1.5B math tokens as well as a dataset of step-by-step math solutions. For this paper, we wanted to explore few-shot in-context learning methods, which is typically used in real-world applications with API-based LLMs.
Self-correction
Pan et al. (2023) present a plethora of self-correction methods in recent literature. While their list includes training-time correction strategies such as RLHF (Ouyang et al., 2022) and self-improve (Huang et al., 2022), our backtracking method falls into the category of post-hoc correction, where the correction process is applied to outputs that have already been generated.
Our paper focuses on correction of logical and reasoning errors, rather than stylistic or qualitative improvements. Previous post-hoc correction methods that are applied to reasoning errors include Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023), both of which cause performance deterioration when the oracle label is not used (Huang et al., 2023). Other methods such as Self-Refine (Madaan et al., 2023) and iterative refinement (Chen et al., 2023) focus on qualitative or stylistic improvements rather than correcting logical errors.
@@ 6 Conclusion
In this paper, we describe and release our dataset BIG-Bench Mistake for mistake finding, and propose a backtracking method to correct logical errors in CoT style traces. We show that LLMs generally struggle with finding logical errors without external feedback, but argue that this feedback can come from a reward model instead. Finally, we demonstrate the effectiveness of backtracking, both with gold standard labels as well as with simulated reward models at lower levels of accuracy.
@@
Another limitation is that our paper does not evaluate backtracking on the original datasets on BIG-Bench, only showing results on the limited set that we sampled in a skewed manner, in order to maximise the value of the human annotators’ time. We leave the full evaluation to future work.
Appendix A Implementational details
@@
We use PaLM 2 L (Unicorn) to generate the traces used in BIG-Bench Mistake. All traces are generated at temperature .
@@
Example CoT traces in the prompt is broken up into smaller steps (typically one sentence per step). This is done so that mistake location information is more precise.
@@
For the logical deduction task, we find that the notation used in the original prompt with question marks is often inconsistent. It becomes difficult for annotators to determine whether a question mark is a mistake or not, because the correctness of the question mark is dependent on its interpretation. To minimise such ambiguity, the question mark notation is rewritten into text.
@@
Following Yao et al. (2022), we use the newline as the stop token, which generates one step with every generation call. We algorithmically append “Thought N:” before each step. This allows us to split up steps in a clear and systematic way. We stop generating once an answer is reached, which is detected using the following regex:
@@ Appendix B Annotation
We release our annotation guidelines at https://github.com/WHGTyen/BIG-Bench-Mistake.
@@
Figure 3 contains an example screenshot of the user interface. For every trace, we provide the input question as well as the target answer, with a note to be aware of errors that may occur in correct traces.
@@
| Model |
Direct (trace) |
Direct (step) |
CoT (step) |
@@
| GPT-4-Turbo |
67.74 |
38.24 |
| GPT-4 |
88.24 |
82.35 |
58.82 |
| GPT-3.5-Turbo |
100.00 |
97.06 |
20.59 |
@@
| GPT-4-Turbo |
90.00 |
77.50 |
| GPT-4 |
82.50 |
82.50 |
80.00 |
| GPT-3.5-Turbo |
67.50 |
0.00 |
0.00 |
@@
| GPT-4-Turbo |
100.00 |
83.33 |
| GPT-4 |
100.00 |
100.00 |
0.00 |
| GPT-3.5-Turbo |
100.00 |
50.00 |
100.00 |
@@
| GPT-4-Turbo |
57.69 |
40.32 |
| GPT-4 |
53.23 |
46.77 |
27.42 |
| GPT-3.5-Turbo |
96.77 |
79.03 |
58.06 |
@@
| GPT-4-Turbo |
96.42 |
30.00 |
| GPT-4 |
98.41 |
78.57 |
13.79 |
| GPT-3.5-Turbo |
95.74 |
4.76 |
0.00 |
| Model |
Direct (trace) |
Direct (step) |
CoT (step) |
@@
| GPT-4-Turbo |
32.71 |
32.33 |
| GPT-4 |
28.20 |
39.47 |
30.83 |
| GPT-3.5-Turbo |
0.00 |
4.51 |
15.04 |
@@
| GPT-4-Turbo |
31.54 |
59.23 |
| GPT-4 |
59.14 |
62.69 |
92.31 |
| GPT-3.5-Turbo |
1.17 |
1.92 |
21.92 |
@@
| GPT-4-Turbo |
20.81 |
74.83 |
| GPT-4 |
39.46 |
67.01 |
10.54 |
| GPT-3.5-Turbo |
0.00 |
24.83 |
7.82 |
@@
| GPT-4-Turbo |
34.27 |
44.12 |
| GPT-4 |
41.60 |
41.60 |
44.54 |
| GPT-3.5-Turbo |
0.00 |
12.18 |
16.81 |
@@
| GPT-4-Turbo |
6.99 |
28.46 |
| GPT-4 |
7.37 |
40.81 |
43.91 |
| GPT-3.5-Turbo |
1.28 |
6.05 |
2.08 |

### v3

@@ 4 Can LLMs correct reasoning mistakes in CoT traces?
We then determine the mistake location in this trace, either from oracle labels (in this section) or with a classifier (in section 5).
If there are no mistakes, we move onto the next trace. If there is a mistake (e.g. at Thought 4 in the example trace in Table 1), we prompt the model again for the same step but at temperature . We use same prompt and the partial trace containing all steps up to but not including the mistake step (e.g. up to Thought 3, prompting for Thought 4).
In our experiments, we found that (c) often produced steps that are identical to the original. We therefore repeat (c) until a different step is generated (or up to a fixed number, whichever is less). For this paper, we use 8 as the maximum number of re-generations; the effects of varying this number is left for future investigation. To reduce computational cost, we generate 8 outputs simultaneously but only select one for backtracking.
@@
This backtracking method is designed to be a very simple baseline, with no specific prompt text or phrasing, and without relying on generating a large number of alternatives. For our experimental results below, we specifically use the same model (PaLM 2 Unicorn) to correct the traces it originally generated, to test its ability to self-correct.
With mistake location | With random location | Avg. num. of steps | |||
@@
Word sorting | -11.11 | +23.53 | -15.56 | +11.76 | 11.7 |
Tracking shuffled objects | -6.67 | +43.92 | -6.67 | +20.39 | 5.4 |
Logical deduction | -11.43 | +36.86 | -13.33 | +21.57 | 8.3 |
Multistep arithmetic | -0.00 | +18.04 | -8.89 | +10.59 | 5.0 |
Dyck languages | -6.82 | +18.06 | -15.91 | +5.16 | 24.5 |
4.1 Results
The results are shown in Table 6. To show that performance increases are not due to randomly resampling outputs, we compare our results to a random baseline, where a mistake location888As described above, the mistake location can be either the number representing the step, or that there are no mistakes. If there are no mistakes, we do not use backtracking and simply use the original trace. is randomly selected for each trace and we perform backtracking based on the random location.
Note that Table 6 separates results into numbers for the correct set and the incorrect set, referring to whether the original trace was correctans or not. This gives a clearer picture than the overall accuracyans, which would be skewed by the proportion of traces that were originally correctans (15%) and incorrectans (85%).
Scores represent the absolute differences in accuracyans. We perform backtracking on both correctans and incorrectans traces, as long as there is a mistake in one of the steps.
@@ 4.1 Results
refers to differences in accuracyans on the set of traces whose original answer was correctans. Note that we take losses here because, despite the correct answer, there is a logical mistake in one of the steps. Therefore, the answer may change to an incorrect one when we backtrack.
@@
is the same but for incorrectans traces, so the answers may have been corrected, hence increasing accuracyans.
@@
For example, for the word sorting task, 11.11% of traces that were originally correctans became incorrectans, while 23.53% of traces that were originally incorrectans became correctans.
4.2 Discussion
@@
The scores show that the gains from correcting incorrectans traces are larger than losses from changing originally correct answers. Additionally, while the random baseline also obtained improvements, they are considerably smaller than if the true mistake location was used. Note that tasks involving fewer steps are more likely to improve performance in the random baseline, as the true mistake location is more likely to be identified.
Our results show that, with mistake location information available, LLMs can correct their own outputs and improve overall downstream performance. This suggests that the main bottleneck in self-correction methods is the identification of mistakes, rather than the correction process. This bottleneck can be overcome by using ground truth feedback (as in Reflexion (Shinn et al., 2023) or RCI (Kim et al., 2023)), or by training a classifier (see section 5).
@@
While our numbers do show that our gains are higher than our losses, it should be noted that changes in the overall accuracy depends on the original accuracy achieved on the task. For example, if the original accuracy on the tracking shuffled objects task was 50%, the new accuracy would be 68.6%. On the other hand, if the accuracy was 99%, the new accuracy would drop to 92.8%. As our dataset is highly skewed and only contains 45 correctans traces per task, we leave to future work a more comprehensive assessment of backtracking, as well as the development of more sophisticated ways to incorporate mistake location information into the self-correction loop.
5 Obtaining mistake location information with a trained classifier
As shown in section 4, if mistake location information is available, LLMs can correct their own CoT traces and boost downstream performance. However, these experimental results are based on oracle labels, which are typically not available in downstream tasks.
One possible solution is to obtain mistake location information from a smaller, trained classifier. If training data is available, one might ask why this approach is preferable to simply fine-tuning the larger, generator model. The reasons are:
Training a small classifier is far more efficient in terms of computing resources and available data.
Once the classifier is trained, it can be used with any LLM as the generator and be updated independently. This can be especially helpful with API-based LLMs that cannot be fine-tuned.
The process of mistake finding is more interpretable than updating the weights of the generator model directly. It clearly pinpoints the location at which an error occurs, which can help the debugging process and allow faster development and iterations of models.
In this section, we seek to answer two questions in the following subsections:
5.1: What mistake-finding accuracy is required for backtracking to be effective?
A trained classifier is unlikely to reach 100% mistake-finding accuracy. If backtracking is only effective when mistake location is 100% accurate, we would not be able to replace oracle labels with a trained classifier.
5.2: Is it possible to improve on results in section 3 without in-domain training data?
Sufficient in-domain training data typically guarantees a performance boost, but can be hard to obtain. We investigate whether mistake-finding in reasoning traces is transferable across tasks. If so, one can use BIG-Bench Mistake or similar datasets to fine-tune a mistake-finding classifier for other tasks.
5.1 Minimum mistake finding accuracy
To explore what level of mistake-finding accuracy is needed, we simulate classifiers at different levels of accuracy and run backtracking for each level. We use accuracymis to refer to the mistake-finding accuracy classifier, to differentiate from downstream task accuracyans.
@@ 5.1 Minimum mistake finding accuracy
For a given classifier at % accuracyclf, we use the mistake location from BIG-Bench Mistake % of the time. For the remaining %, we sample a mistake location randomly. To mimic the behaviour of a typical classifier, mistake locations are sampled to match the distribution found in the dataset. We also ensure that the sampled location does not match the correct location.
@@
Results are shown in Figure 3. We can see that the losses in accuracy begins to plateau at 65%. In fact, for most tasks, accuracy is already larger than accuracy at around 60-70% accuracyRM. This demonstrates that while higher accuracies produce better results, backtracking is still effective even without gold standard mistake location labels.
5.2 Training a classifier on out-of-domain data
@@
We test whether mistake-finding can benefit from a dedicated classifier trained on out-of-distribution tasks. We fine-tune PaLM 2 Otter, a model much smaller than PaLM 2 Unicorn, with our BIG-Bench Mistake data for 20k steps and choose the checkpoint with the best validation results. For each task, we hold out the in-domain data for evaluation while training the classifier on the other 4 tasks.
Held-out task | Trained classifier accuracymis (Otter) | 3-shot prompting accuracymis (Unicorn) | Difference |
Word sorting | 22.33 | 11.67 | +11.66 |
Tracking shuffled objects | 37.67 | 18.00 | +19.67 |
Logical deduction | 6.00 | 6.67 | -0.67 |
Multi-step arithmetic | 26.00 | 22.00 | +4.00 |
Dyck languages | 33.57 | 10.98 | +22.59 |
@@
We show the relative improvements and losses in Table 7 vs. a 3-shot baseline on PaLM 2 Unicorn (scores from section 3). We see gains for 4 out of 5 of the tasks. Note the classifiers we train are significantly smaller than our inference model, and is trained on out-of-domain data. This suggests that it may be possible to train classifiers to assist in backtracking, and that these classifiers do not have to be large. Further, such a classifier can work on out-of-distribution mistakes.
Despite this, the performance of our classifiers do not meet the threshold required for effective backtracking, as demonstrated in subsection 5.1. We believe more data may be necessary to improve results across the board on all tasks. We leave to future work the collection of this larger dataset and a more rigorous investigation of the trade-offs of model size vs. performance of the classifier.
@@
We also leave for future investigation the effect of backtracking iteratively with a classifier: for example, the generator model may make another mistake after backtracking for the first time, which can then be identified and corrected again.
@@ 6 Related work
6.1 Datasets
To our knowledge, the only publicly available dataset containing mistake annotations in LLM outputs is PRM800K (Lightman et al., 2023), which is a dataset of solutions to Olympiad-level math questions. Our dataset BIG-Bench Mistake covers a wider range of tasks to explore the reasoning capabilities of LLMs more thoroughly. Additionally, our dataset explores the more general case of prompting API-based LLMs, whereas PRM800K uses a generator LLM that was heavily fine-tuned for math-specific problems.
6.2 Self-correction
The term self-correction can be applied to a wide variety of techniques. Pan et al. (2023) present a survey of self-correction methods in recent literature. Broadly, proposed self-correction methods can vary in the following dimensions:
Source of feedback
Some techniques rely on external feedback inherent to the task such as code execution errors (Yasunaga and Liang, 2020; Chen et al., 2023a, 2024b), or feedback from humans (e.g. Chen et al., 2024a; Yuan et al., 2024). Some explicitly train a model to produce feedback (e.g. Madaan et al., 2021; Welleck et al., 2023; Bai et al., 2022b; Lee et al., 2023; Paul et al., 2023; Ouyang et al., 2022; Bai et al., 2022a; Ganguli et al., 2023), while others rely on prompting only (e.g. Shinn et al., 2023; Miao et al., 2024). It has been demonstrated that prompting-only setups can work well for stylistic or qualitative improvements (e.g. Madaan et al., 2023; Chen et al., 2023b), but would require external feedback for reasoning tasks (Huang et al., 2023; Shinn et al., 2023; Kim et al., 2023; Madaan et al., 2023).
Time of correction
Feedback can be incorporated at various points during the self-correction process. Some methods do so by updating weights during training time (e.g. Ouyang et al., 2022; Bai et al., 2022a; Ganguli et al., 2023); some do so during generation time (e.g. Weng et al., 2023; Dalvi Mishra et al., 2022; Xie et al., 2023); others apply correction to output that has already been generated (e.g. Saunders et al., 2022; Kim et al., 2023; Shinn et al., 2023). Our method falls into the final category (post-hoc correction), as it involves identifying an incorrect step in a complete CoT trace; however, it is also possible to apply a mistake-finding classifier at every step during generation.
@@ 7 Conclusion
In this paper, we investigate LLMs’ ability to find mistakes and correct outputs. We find that LLMs generally struggle to find mistakes, but, when given mistake location information, are able to correct outputs to boost performance. We therefore hypothesise that mistake finding is an important bottleneck preventing self-corrections strategies from performing well on reasoning tasks.
We show initial evidence that a dedicated classifier for mistake finding can overcome this bottleneck. We train a small, baseline classifier on out-of-domain data and demonstrate improvement over few-shot prompting results. While the classifier does not reach the threshold required for effective backtracking, our results show that it is possible to improve on mistake-finding accuracy using standard machine learning techniques. We leave the development of more sophisticated methods to future work, and release our dataset BIG-Bench Mistake to encourage this direction of research.
@@
Another limitation is that our paper does not experiment with backtracking on the original datasets on BIG-Bench, only showing results on the limited set that we sampled in a skewed manner, in order to maximise the value of the human annotators’ time. We leave the full evaluation to future work as this is beyond the scope of this paper, which is intended as a proof-of-concept to show the importance of mistake-finding.
Acknowledgements
We thank Vicky Zayats and Sian Gooding for their invaluable feedback, as well as members of the Google Research and Google DeepMind teams who gave continued support throughout the project.
Appendix A Dataset details
Our dataset, BIG-Bench Mistake, is available at https://github.com/WHGTyen/BIG-Bench-Mistake under the Apache License 2.0. The five tasks used in our dataset are based on BIG-Bench (Srivastava et al., 2022), also released under the Apache License 2.0. All five tasks are in the English language.
@@ Appendix A Dataset details
We use PaLM 2 (Unicorn) to generate the traces used in BIG-Bench Mistake. All traces are generated at temperature .
@@
Example CoT traces in the prompt are broken up into smaller steps (typically one sentence per step). This is done so that mistake location information is more precise.
@@
For the logical deduction task, we find that the notation used in the original prompt with question marks is often inconsistent. It becomes difficult for annotators to determine whether a question mark is a mistake or not, because the correctness of the question mark is dependent on its interpretation. To minimise such ambiguity, we replace the question mark notation with text descriptions of the objects.
@@
Following Yao et al. (2022), we use the newline as the stop token, thereby generating one step with every generation call. We algorithmically append “Thought N:” before each step. This allows us to split up steps in a clear and systematic way. We stop generating once an answer is reached, which is detected using the following regex:
@@ Appendix B Annotation
We release our annotation guidelines at https://github.com/WHGTyen/BIG-Bench-Mistake. Our annotators are recruited via our institution and contracted at the market rate in their country of residence.
@@
Figure 4 contains an example screenshot of the user interface. For every trace, we provide the input question as well as the target answer, with a note to be aware of errors that may occur in correctans traces.
@@
Appendix C Training mistake-finding classifiers
To train our mistake-finding classifiers (see subsection 5.2), we fine-tune PaLM 2 Otter on 4 of our 5 tasks, holding out one task for evaluation. This is done for each of our 5 tasks.
All 5 models are fine-tuned for 20k steps with a batch size of 32. The learning rate is with a linear ramp and cosine decay. After 20k steps, we select the checkpoint with the best validation results. The number of steps trained for each model are shown in Table 8.
Held-out task | Training steps |
Word sorting | 6800 |
Tracking shuffled objects | 8000 |
Logical deduction | 9000 |
Multi-step arithmetic | 10000 |
Dyck languages | 10000 |
All models are trained as a binary classifier on whether a CoT step is correct, given the task and previous steps. Due to the limited data, we include in training the CoT steps that occur after the first mistake step. These steps are considered incorrect for the purposes of training (despite not being human-annotated as such).
Appendix D User interface
Appendix E Benchmark scores
Model | Direct (trace) | Direct (step) | CoT (step) |
@@ Appendix E Benchmark scores
GPT-4-Turbo | 67.74 | 38.24 | – |
GPT-4 | 88.24 | 82.35 | 58.82 |
GPT-3.5-Turbo | 100.00 | 97.06 | 20.59 |
Gemini Pro | 44.12 | – | – |
PaLM 2 Unicorn | 100.00 | 73.53 | 35.29 |
@@
GPT-4-Turbo | 90.00 | 77.50 | – |
GPT-4 | 82.50 | 82.50 | 80.00 |
GPT-3.5-Turbo | 67.50 | 0.00 | 0.00 |
Gemini Pro | 12.50 | – | – |
PaLM 2 Unicorn | 100.00 | 85.00 | 47.50 |
@@
GPT-4-Turbo | 100.00 | 83.33 | – |
GPT-4 | 100.00 | 100.00 | 0.00 |
GPT-3.5-Turbo | 100.00 | 50.00 | 100.00 |
Gemini Pro | 33.33 | – | – |
PaLM 2 Unicorn | 100.00 | 100.00 | 50.00 |
@@
GPT-4-Turbo | 57.69 | 40.32 | – |
GPT-4 | 53.23 | 46.77 | 27.42 |
GPT-3.5-Turbo | 96.77 | 79.03 | 58.06 |
Gemini Pro | 83.87 | – | – |
PaLM 2 Unicorn | 83.87 | 93.55 | 29.03 |
@@
GPT-4-Turbo | 96.42 | 30.00 | – |
GPT-4 | 98.41 | 78.57 | 13.79 |
GPT-3.5-Turbo | 95.74 | 4.76 | 0.00 |
Gemini Pro | 0.00 | – | – |
PaLM 2 Unicorn | 100.00 | 80.95 | 19.05 |
Model | Direct (trace) | Direct (step) | CoT (step) |
@@
GPT-4-Turbo | 32.71 | 32.33 | – |
GPT-4 | 28.20 | 39.47 | 30.83 |
GPT-3.5-Turbo | 0.00 | 4.51 | 15.04 |
Gemini Pro | 6.39 | – | – |
PaLM 2 Unicorn | 0.38 | 9.02 | 11.28 |
@@
GPT-4-Turbo | 31.54 | 59.23 | – |
GPT-4 | 59.14 | 62.69 | 92.31 |
GPT-3.5-Turbo | 1.17 | 1.92 | 21.92 |
Gemini Pro | 41.54 | – | – |
PaLM 2 Unicorn | 5.38 | 19.23 | 56.92 |
@@
GPT-4-Turbo | 20.81 | 74.83 | – |
GPT-4 | 39.46 | 67.01 | 10.54 |
GPT-3.5-Turbo | 0.00 | 24.83 | 7.82 |
Gemini Pro | 8.16 | – | – |
PaLM 2 Unicorn | 4.76 | 36.73 | 11.22 |
@@
GPT-4-Turbo | 34.27 | 44.12 | – |
GPT-4 | 41.60 | 41.60 | 44.54 |
GPT-3.5-Turbo | 0.00 | 12.18 | 16.81 |
Gemini Pro | 5.46 | – | – |
PaLM 2 Unicorn | 5.88 | 2.94 | 22.27 |
@@
GPT-4-Turbo | 6.99 | 28.46 | – |
GPT-4 | 7.37 | 40.81 | 43.91 |
GPT-3.5-Turbo | 1.28 | 6.05 | 2.08 |
Gemini Pro | 2.25 | – | – |
PaLM 2 Unicorn | 0.38 | 6.43 | 17.77 |